# Zhihui Zhu

### v1

@@ 4 Proofs of Main Results
In this section, we sketch the proofs of our main theorems in Section 3. The complete proofs can be found in Sections B and C of the appendix. For ease of exposition, let
$\alpha:=\frac{d}{m\epsilon^{2}},\quad\alpha_{k}:=\frac{d}{m_{k}\epsilon^{2}},\quad\forall k\in[K].$ (11)
@@ 4.1 Analysis of Optimality Conditions
Our goal in this subsection is to characterize the local and global optima of Problem (5). Towards this goal, we first provide an upper bound on the objective function $F$ in Problem (5). In particular, this upper bound is tight when the blocks $\{\bm{Z}_{k}\}_{k=1}^{K}$ are orthogonal to each other. This result is a direct consequence of [4, Lemma 10].
@@
Consider the setting of Theorem 1. It holds that $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\cal Z$ if and only if each $\bm{Z}_{k}$ admits the following singular value decomposition
@@
$\overline{\sigma}_{k}=\left(\frac{\eta_{k}+\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2},\quad\underline{\sigma}_{k}=\left(\frac{\eta_{k}-\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2}.$ (16)
@@ 4.2 Analysis of Optimization Landscape
Moreover, according to Proposition 2, we further divide $\cal Z$ into two disjoint sets $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Here,
@@
Consider the setting of Theorem 2. Suppose that $\bm{Z}\in\cal Z$. Then, the following statements hold:
@@
Next, we proceed to the second step to show that any point belonging to $\mathcal{Z}^{c}$ is a strict saddle point. It suffices to find a direction $\bm{D}\in\mathbb{R}^{d\times m}$ such that $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0$ for each $\bm{Z}\in\mathcal{Z}^{c}$ according to Definition 2.
@@
Consider the setting of Theorem 2. If $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point and there exists $1\leq k\neq l\leq K$ such that $\bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0}$, i.e., $\bm{Z}\in\mathcal{Z}^{c}$, then $\bm{Z}$ is a strict saddle point.
With the above preparations that characterize all the critical points, we can prove Theorem 1 and Theorem 2. We refer the reader to Appendix D for the detailed proof.
@@ 5 Experimental Results
In this section, we first conduct numerical experiments on synthetic data in Section 5.1 to validate our theoretical results, and then on real-world data sets using deep neural networks in Section 5.2 to further support our theory. All codes are implemented in Python mainly using NumPy and PyTorch. All of our experiments are executed on a computing server equipped with NVIDIA A40 GPUs. Due to space limitations, we defer some implementation details and additional experimental results to Appendix E.
@@ 5.1 Validation of Theory for Solving Problem (5)
Verification of Theorem 1.
In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=30,m_{2}=70,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. Then, one can verify that $\lambda$ satisfies (6). For the solution $\bm{Z}$ returned by GD, we first plot the heatmap of the cosine similarity between pairwise columns of $\bm{Z}$ in Figure 2(a). We observe that the features from different classes are orthogonal to each other, while the features from the same class are correlated. Next, we compute the singular values of $\bm{Z}_{k}$ via singular value decomposition (SVD) and plot the singular values using blue dots for each $k\in[K]$ in Figure 2(b). According to the closed-form solution (7) in Theorem 1, we also plot the theoretical bound of singular values in red in Figure 2(b). One can observe that the number of singular values of each block is respectively $24,23,27,26$, summing up to $100$, and the red line perfectly matches the blue dots. These results all provide strong support for Theorem 1.
Verification of Theorem 2.
In this experiment, we maintain the same setting as above, except that the number of samples in each class is equal. We first fix $m=200$ and vary $d\in\{40,80,120\}$, and then fix $d=50$ and vary $d\in\{100,200,400\}$ to run GD. We plot the distances between function values of the iterates to the optimal value, which is computed according to (7) in Theorem 1, against the iteration numbers in Figure 3. We observe that GD with random initialization converges to an optimal solution at a linear rate. This indicates that the MCR2 has a benign global landscape, which supports Theorem 2.
@@
Figure 3: Convergence performance of GD for solving the regularized MCR2 problem. Here, the $x$-axis is number of iterations (also denoted by $t$), and $y$-axis is the function value gap $F^{t}-F^{*}$, where $F^{t}=F(\bm{Z}^{t})$ denotes the function value at the $t$-th iterate $\bm{Z}^{t}$ generated by GD, and $F^{*}$ is the optimal value of Problem (5) computed according to (7) in Theorem 1.
@@ 5.2 Training Deep Networks Using Regularized MCR2
In this subsection, we conduct numerical experiments on the image datasets MNIST [23] and CIFAR-10 [22] to provide evidence that our theory also applies to deep networks. More specifically, we employ a multi-layer perceptron network with ReLU activation as the feature mapping $\bm{z}=f_{\bm{\Theta}}(\bm{x})$ with output dimension 32 for MNIST and 128 for CIFAR-10. Then, we train the network parameters $\bm{\Theta}$ via Adam [20] by optimizing Problem (5).
@@
In the experiments, we randomly sample a balanced subset with $K$ classes and $m$ samples from MNIST or CIFAR-10, where each class has the same number of samples. We set $\lambda=0.001$ and $\epsilon=0.5$. For different subsets with corresponding values of $(m,K)$, we run experiments and report the function value $\hat{F}$ obtained by training deep networks and the optimal value $F^{*}$ computed using the closed-form solution in Theorem 1 in Table 1. To verify the discriminative nature of the features obtained by training deep networks across different classes, we measure the discrimination between features belonging to different classes by computing the cosine of the principal angle [3] between the class subspaces: $s=\max\left\{\|\bm{U}_{k}^{T}\bm{U}_{l}\|:k\neq l\in[K]\right\}\in[0,1]$, where the columns of $\bm{U}_{k}\in\mathbb{R}^{d\times r_{k}}$ are the right singular vectors corresponding to the top $r_{k}$ singular values of $\bm{Z}_{k}$ defined in (14) and $r_{k}$ is its rank [^6] We estimate the rank of a matrix by rounding its “stable rank” [17]: $r_{k}=\operatorname{round}(\|\bm{Z}_{k}\|_{F}^{2}/\|\bm{Z}_{k}\|^{2})$. for each $k\in[K]$. In particular, when $s$ is smaller, the spaces spanned by each pair $\bm{Z}_{k}$ and $\bm{Z}_{l}$ for $k\neq l$ are closer to being orthogonal to each other. Then, we record the value $s$ in Table 1 in different settings. Moreover, we visualize the pairwise cosine similarities between learned features on MNIST and CIFAR-10 when $(m,K)=(1500,6)$ and $(2500,10)$ in Figure 4.
@@
We observe from Table 1 that the function value returned by training deep networks is extremely close to the global optimal value of Problem (5) and from the value $s$ and Figure 4 that the features from different classes are nearly orthogonal to each other. These observations, together with Theorems 1 and 2, indicate that Problem (5) retains its optimization properties even when $\bm{Z}$ is parameterized by a neural network. Our theoretical analysis of Problem (5) thus illustrates a qualitative picture of training deep networks with the regularized MCR2 objective.
@@ 6 Conclusion
In this work, we provided a complete characterization of the global landscape of the MCR2 objective, a highly nonconcave and nonlinear function used for representation learning. We characterized all critical points, including the local and global optima, of the MCR2 objective, and showed that — surprisingly — it has a benign global optimization landscape. These characterizations provide rigorous justifications for why such an objective can be optimized well using simple algorithms such as gradient-based methods. In particular, we show that even local optima of the objective leads to geometrically meaningful representations. Our experimental results on synthetic and real-world datasets clearly support this new theoretical characterization. With the global landscape clearly revealed, our work paves the way for exploring better optimization strategies, hence better deep neural network architectures, for optimizing the MCR2 objective more efficiently and effectively. For future work, it is natural to extend our analysis to Problem (4) with deep network parameterizations. It is also interesting to study the sparse MCR2 objective, which has led to high-performance transformer-like architectures [57, 58].
@@ Acknowledgements
The organization of the supplementary material is as follows: In Appendix A, we introduce preliminary setups and auxiliary results for studying the MCR2 problem. Then, we prove the technical results concerning the global optimality of Problem (5) in Appendix B and the optimization landscape of Problem (5) in Appendix C, respectively. In Appendix D, we prove the main theorems in Theorem 1 and Theorem 2. Finally, we provide more experimental setups and results in Appendix E.
@@ A.1 Optimality Conditions and Strict Saddle Points
where $\bm{X}:=\bm{I}_{n}+\alpha\bm{Z}\bm{Z}^{T}$ and $\alpha$ is defined in (11). Note that we can compute the gradient and Hessian of $R_{c}(\cdot)$ in (20) using the same approach. Based on the above setup, we define the first-order optimality condition of Problem (5) as follows.
@@
According to Jin et al. [19], Lee et al. [26], we define the strict saddle point, i.e., a critical point that has a direction with strictly positive curvature [^7] Note that Problem (5) is not a minimization problem but a maximization problem., of Problem (5) as follows:
@@
Remark that for the MCR2 problem, strict saddle points include classical saddle points with strictly positive curvature as well as local minimizers.
@@ A.2 Auxiliary Results
We next present the commutative property for the log-determinant function and the upper bound for the coding rate function. We refer the reader to [4, Lemma 8 & Lemma 10] for the detailed proofs. Here, let $\bm{Z}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decompositon of $\bm{Z}\in\mathbb{R}^{d\times m}$, where $r=\mathrm{rank}(\bm{Z})\leq\min\{m,d\}$, $\bm{U}\in\mathcal{O}^{d\times r}$, $\bm{\Sigma}\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $\bm{V}\in\mathcal{O}^{m\times r}$.
@@
$F(\bm{Z}\bm{O})=F(\bm{Z}),\quad\nabla F(\bm{Z}\bm{O})=\nabla F(\bm{Z})\bm{O},\quad\nabla^{2}F(\bm{Z}\bm{O})[\bm{D}\bm{O},\bm{D}\bm{O}]=\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}].$ (29)
@@
[…]
Appendix B Proofs in Section 4.1
B.1 Proof of Lemma 1
Proof of Lemma 1.
It follows from (28) in Lemma 4 that
@@ B.1 Proof of Lemma 1
B.2 Proof of Proposition 2
Proof of Proposition 2.
@@ B.2 Proof of Proposition 2
$\eta_{k}>0,\quad\eta_{k}^{2}-\frac{4\alpha_{k}}{\alpha}\lambda^{2}\geq 0.$
@@
Finally, using $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ and (33), we obtain $\bm{Q}_{k,1}^{T}\bm{Q}_{l,1}=\bm{0}$ for all $1\leq l\neq k\leq K$. These, together with (33), yields (14).
@@
Appendix C Proofs in Section 4.2
C.1 Proof of Proposition 3
Proof of Proposition 3.
For each $\bm{Z}\in\cal Z$, it follows from Lemma 1 that

### v2

@@ 4 Proofs of Main Results
In this section, we sketch the proofs of our main theorems in Section˜3. The complete proofs can be found in Sections B and C of the appendix. For ease of exposition, let
$\alpha:=\frac{d}{m\epsilon^{2}},\qquad\alpha_{k}:=\frac{d}{m_{k}\epsilon^{2}},\qquad\forall k\in[K].$ (11)
@@ 4.1 Analysis of Optimality Conditions
Our goal in this subsection is to characterize the local and global optima of Problem (5). Towards this goal, we first provide an upper bound on the objective function $F$ in Problem (5). In particular, this upper bound is tight when the blocks $\{\bm{Z}_{k}\}_{k=1}^{K}$ are orthogonal to each other. This result is a direct consequence of [chan2022redunet, Lemma 10].
@@
Consider the setting of Theorem˜1. It holds that $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\cal Z$ if and only if each $\bm{Z}_{k}$ admits the following singular value decomposition
@@
$\overline{\sigma}_{k}=\left(\frac{\eta_{k}+\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2},\qquad\underline{\sigma}_{k}=\left(\frac{\eta_{k}-\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2}.$ (16)
@@ 4.2 Analysis of Optimization Landscape
Moreover, according to ˜2, we further divide $\cal Z$ into two disjoint sets $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Here,
@@
Consider the setting of Theorem˜2. Suppose that $\bm{Z}\in\cal Z$. Then, the following statements hold:
@@
Next, we proceed to the second step to show that any point belonging to $\mathcal{Z}^{c}$ is a strict saddle point. It suffices to find a direction $\bm{D}\in\mathbb{R}^{d\times m}$ such that $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0$ for each $\bm{Z}\in\mathcal{Z}^{c}$ according to Definition˜2.
@@
Consider the setting of Theorem˜2. If $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point and there exists $1\leq k\neq l\leq K$ such that $\bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0}$, i.e., $\bm{Z}\in\mathcal{Z}^{c}$, then $\bm{Z}$ is a strict saddle point.
With the above preparations that characterize all the critical points, we can prove Theorem˜1 and Theorem˜2. We refer the reader to Appendix˜D for the detailed proof.
@@ 5 Experimental Results
In this section, we first conduct numerical experiments on synthetic data in Section˜5.1 to validate our theoretical results, and then on real-world data sets using deep neural networks in Section˜5.2 to further support our theory. All codes are implemented in Python mainly using NumPy and PyTorch. All of our experiments are executed on a computing server equipped with NVIDIA A40 GPUs. Due to space limitations, we defer some implementation details and additional experimental results to Appendix˜E.
@@ 5.1 Validation of Theory for Solving Problem (5)
Verification of Theorem˜1.
In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=30,m_{2}=70,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. Then, one can verify that $\lambda$ satisfies (6). For the solution $\bm{Z}$ returned by GD, we first plot the heatmap of the cosine similarity between pairwise columns of $\bm{Z}$ in Figure˜2(a). We observe that the features from different classes are orthogonal to each other, while the features from the same class are correlated. Next, we compute the singular values of $\bm{Z}_{k}$ via singular value decomposition (SVD) and plot the singular values using blue dots for each $k\in[K]$ in Figure˜2(b). According to the closed-form solution (7) in Theorem˜1, we also plot the theoretical bound of singular values in red in Figure˜2(b). One can observe that the number of singular values of each block is respectively $24,23,27,26$, summing up to $100$, and the red line perfectly matches the blue dots. These results all provide strong support for Theorem˜1.
Verification of Theorem˜2.
In this experiment, we maintain the same setting as above, except that the number of samples in each class is equal. We first fix $m=200$ and vary $d\in\{40,80,120\}$, and then fix $d=50$ and vary $d\in\{100,200,400\}$ to run GD. We plot the distances between function values of the iterates to the optimal value, which is computed according to (7) in Theorem˜1, against the iteration numbers in Figure˜3. We observe that GD with random initialization converges to an optimal solution at a linear rate. This indicates that the MCR2 has a benign global landscape, which supports Theorem˜2.
@@
Figure 3: Convergence performance of GD for solving the regularized MCR2 problem. Here, the $x$-axis is number of iterations (also denoted by $t$), and $y$-axis is the function value gap $F^{t}-F^{*}$, where $F^{t}=F(\bm{Z}^{t})$ denotes the function value at the $t$-th iterate $\bm{Z}^{t}$ generated by GD, and $F^{*}$ is the optimal value of Problem (5) computed according to (7) in Theorem˜1.
@@ 5.2 Training Deep Networks Using Regularized MCR2
In this subsection, we conduct numerical experiments on the image datasets MNIST [lecun1998gradient] and CIFAR-10 [krizhevsky2009learning] to provide evidence that our theory also applies to deep networks. More specifically, we employ a multi-layer perceptron network with ReLU activation as the feature mapping $\bm{z}=f_{\bm{\Theta}}(\bm{x})$ with output dimension 32 for MNIST and 128 for CIFAR-10. Then, we train the network parameters $\bm{\Theta}$ via Adam [kingma2014adam] by optimizing Problem (5).
@@
In the experiments, we randomly sample a balanced subset with $K$ classes and $m$ samples from MNIST or CIFAR-10, where each class has the same number of samples. We set $\lambda=0.001$ and $\epsilon=0.5$. For different subsets with corresponding values of $(m,K)$, we run experiments and report the function value $\hat{F}$ obtained by training deep networks and the optimal value $F^{*}$ computed using the closed-form solution in Theorem˜1 in Table˜1. To verify the discriminative nature of the features obtained by training deep networks across different classes, we measure the discrimination between features belonging to different classes by computing the cosine of the principal angle [bjorck1973numerical] between the class subspaces: $s=\max\left\{\|\bm{U}_{k}^{T}\bm{U}_{l}\|:k\neq l\in[K]\right\}\in[0,1]$, where the columns of $\bm{U}_{k}\in\mathbb{R}^{d\times r_{k}}$ are the right singular vectors corresponding to the top $r_{k}$ singular values of $\bm{Z}_{k}$ defined in (14) and $r_{k}$ is its rank [^6] We estimate the rank of a matrix by rounding its “stable rank” [horn2012matrix]: $r_{k}=\operatorname{round}(\|\bm{Z}_{k}\|_{F}^{2}/\|\bm{Z}_{k}\|^{2})$. for each $k\in[K]$. In particular, when $s$ is smaller, the spaces spanned by each pair $\bm{Z}_{k}$ and $\bm{Z}_{l}$ for $k\neq l$ are closer to being orthogonal to each other. Then, we record the value $s$ in Table˜1 in different settings. Moreover, we visualize the pairwise cosine similarities between learned features on MNIST and CIFAR-10 when $(m,K)=(1500,6)$ and $(2500,10)$ in Figure˜4.
@@
We observe from Table˜1 that the function value returned by training deep networks is extremely close to the global optimal value of Problem (5) and from the value $s$ and Figure˜4 that the features from different classes are nearly orthogonal to each other. These observations, together with Theorems˜1 and 2, indicate that Problem (5) retains its optimization properties even when $\bm{Z}$ is parameterized by a neural network. Our theoretical analysis of Problem (5) thus illustrates a qualitative picture of training deep networks with the regularized MCR2 objective.
@@ 6 Conclusion
In this work, we provided a complete characterization of the global landscape of the MCR2 objective, a highly nonconcave and nonlinear function used for representation learning. We characterized all critical points, including the local and global optima, of the MCR2 objective, and showed that — surprisingly — it has a benign global optimization landscape. These characterizations provide rigorous justifications for why such an objective can be optimized well using simple algorithms such as gradient-based methods. In particular, we show that even local optima of the objective leads to geometrically meaningful representations. Our experimental results on synthetic and real-world datasets clearly support this new theoretical characterization. With the global landscape clearly revealed, our work paves the way for exploring better optimization strategies, hence better deep neural network architectures, for optimizing the MCR2 objective more efficiently and effectively. For future work, it is natural to extend our analysis to Problem (4) with deep network parameterizations. It is also interesting to study the sparse MCR2 objective, which has led to high-performance transformer-like architectures [yu2023sparse, yu2023white].
@@ Acknowledgements
The organization of the supplementary material is as follows: In Appendix˜A, we introduce preliminary setups and auxiliary results for studying the MCR2 problem. Then, we prove the technical results concerning the global optimality of Problem (5) in Appendix˜B and the optimization landscape of Problem (5) in Appendix˜C, respectively. In Appendix˜D, we prove the main theorems in Theorem˜1 and Theorem˜2. Finally, we provide more experimental setups and results in Appendix˜E.
@@ A.1 Optimality Conditions and Strict Saddle Points
where $\bm{X}:=\bm{I}_{d}+\alpha\bm{Z}\bm{Z}^{T}$ and $\alpha$ is defined in (11). Note that we can compute the gradient and Hessian of $R_{c}(\cdot)$ in (20) using the same approach. Based on the above setup, we define the first-order optimality condition of Problem (5) as follows.
@@
According to jin2017escape, lee2019first, we define the strict saddle point, i.e., a critical point that has a direction with strictly positive curvature [^7] Note that Problem (5) is not a minimization problem but a maximization problem., of Problem (5) as follows:
@@
Remark that for the MCR2 problem, strict saddle points include saddle points with strictly positive curvature as well as local maximizers.
@@ A.2 Auxiliary Results
We next present the commutative property for the log-determinant function and the upper bound for the coding rate function. We refer the reader to [chan2022redunet, Lemma 8 & Lemma 10] for the detailed proofs. Here, let $\bm{Z}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decompositon of $\bm{Z}\in\mathbb{R}^{d\times m}$, where $r=\mathrm{rank}(\bm{Z})\leq\min\{m,d\}$, $\bm{U}\in\mathcal{O}^{d\times r}$, $\bm{\Sigma}\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $\bm{V}\in\mathcal{O}^{m\times r}$.
@@
$F(\bm{Z}\bm{O})=F(\bm{Z}),\qquad\nabla F(\bm{Z}\bm{O})=\nabla F(\bm{Z})\bm{O},\qquad\nabla^{2}F(\bm{Z}\bm{O})[\bm{D}\bm{O},\bm{D}\bm{O}]=\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}].$ (29)
Proof of Lemma˜5.
[…]
Appendix B Proofs in Section˜4.1
B.1 Proof of Lemma˜1
@@
It follows from (28) in Lemma˜4 that
@@
B.2 Proof of ˜2
Proof of ˜2.
@@ B.2 Proof of ˜2
$\eta_{k}>0,\qquad\eta_{k}^{2}-\frac{4\alpha_{k}}{\alpha}\lambda^{2}\geq 0.$
@@
Finally, using $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ and (33), we obtain $\bm{P}_{k,1}^{T}\bm{P}_{l,1}=\bm{0}$ for all $1\leq l\neq k\leq K$. These, together with (33), yields (14).
@@
Appendix C Proofs in Section˜4.2
C.1 Proof of ˜3
Proof of ˜3.
For each $\bm{Z}\in\cal Z$, it follows from Lemma˜1 that